### v1

# IMAGHarmony: Controllable Image Editing with Consistent Object Quantity and Layout

Abstract
Recent diffusion models have advanced image editing by enhancing visual quality and control, supporting broad applications across creative and personalized domains. However, current image editing largely overlooks multi-object scenarios, where precise control over object categories, counts, and spatial layouts remains a significant challenge. To address this, we introduce a new task, quantity-and-layout consistent image editing (QL-Edit), which aims to enable fine-grained control of object quantity and spatial structure in complex scenes. We further propose IMAGHarmony, a structure-aware framework that incorporates harmony-aware attention (HA) to integrate multimodal semantics, explicitly modeling object counts and layouts to enhance editing accuracy and structural consistency. In addition, we observe that diffusion models are susceptible to initial noise and exhibit strong preferences for specific noise patterns. Motivated by this, we present a preference-guided noise selection (PNS) strategy that chooses semantically aligned initial noise samples based on vision-language matching, thereby improving generation stability and layout consistency in multi-object editing. To support evaluation, we construct HarmonyBench, a comprehensive benchmark covering diverse quantity and layout control scenarios. Extensive experiments demonstrate that IMAGHarmony consistently outperforms state-of-the-art methods in structural alignment and semantic accuracy. The code and model are available at https://github.com/muzishen/IMAGHarmony.
1 Introduction
Recent advances in image generation [1, 20, 35, 30, 22, 27] are largely driven by diffusion models [7, 48, 50, 67], which outperform traditional methods such as GANs [39, 37] and VAEs [53, 8] across diverse applications. Building on this progress, prompt-driven image editing [41, 42, 32] allows users to flexibly modify existing images while preserving visual fidelity and semantic alignment. These advances have enabled broad adoption in creative, commercial, and personalized content creation.
Most prompt-driven image editing methods [68, 3, 32, 54, 29, 31] fall into two main categories: mask-free [68, 3, 57, 17, 20] and mask-based approaches [2, 69, 65, 71, 49]. Mask-free approaches [68, 3, 57] rely only on textual prompts to steer the diffusion model, producing global changes without explicit spatial constraints. Mask-based approaches [2, 69, 65], in turn, confine the edit to user-defined masks, enabling fine-grained updates space between text and reference image. As shown in Figure 1 (a), both paradigms perform well when editing a few-object (e.g., one or two). However, they [41, 29, 42, 32] often fail in scenes requiring consistent control over more than two objects, particularly in terms of object counts and spatial layouts from Figure 1 (b). In addition, multi-object editing frequently results in incorrect styles (e.g., Cyberpunk Style) and disorganized backgrounds (e.g., Flowers in Background). Two primary factors contribute to this limitation: (i) The weak counting capability of diffusion models, where text prompts often fail to convey or retain the desired object quantity; (ii) The inability to distinguish multiple entities from object parts within masked regions, hindering accurate structural and semantic modeling.
Figure 1: Editing results on (a) few-object and (b) multi-object scenes. Existing methods struggle to preserve the count, layout, and semantics in multi-object cases, while ours ensures consistent and faithful edits.
Furthermore, recent studies [14, 26, 35, 63, 38] have shown that diffusion models are highly susceptible to the choice of initial noise in generation and editing tasks. Since the denoising trajectory [56, 66, 55] is conditioned on the starting noise, different seeds induce distinct attention pathways and layout biases during early stages, leading to substantial variations in object count, spatial arrangement, and visual semantics. For example, InitNO [14] identifies suboptimal noise as a significant cause of prompt-image misalignment and proposes an attention-based strategy to guide noise toward semantically aligned outputs. Building on these insights [64, 63, 36], we observe that initial noise is equally critical in multi-object editing, particularly in determining object count and spatial arrangement. Even under identical prompts, different seeds can produce structurally divergent outputs: some yield clean and well-distributed layouts, while others result in missing, overlapping, or misaligned objects. Therefore, identifying effective noise seeds during inference is crucial for ensuring reliable layout and object consistency in multi-object editing.
To address the above challenges, we present a new task: quantity-and-layout consistent image editing (QL-Edit). Given a reference image and a textual instruction, the objective is to edit the image such that the output not only aligns with the desired semantics but also strictly conforms to the specified object count and spatial arrangement. To this end, we propose IMAGHarmony, a consistency-aware framework that integrates two key components: a harmony-aware attention (HA) module and a preference-guided noise selection (PNS) strategy. Specifically, the HA module addresses the limitations of existing models in quantity and layout modeling by explicitly encoding object count from auxiliary textual features and implicitly capturing spatial relations from visual features. These representations are fused and injected into the denoising network to guide layout-consistent editing. To mitigate the sensitivity of diffusion models to initial noise, the PNS strategy generates multiple candidate seeds during training. At inference, it performs a few-steps denoising for each seed, evaluates semantic alignment using a frozen vision-language model, and selects the best-performing seed to initialize the complete denoising process, thereby improving stability and structural fidelity. To support evaluation, we further construct HarmonyBench, a comprehensive benchmark tailored for QL-Edit. It includes a wide range of multi-object editing scenarios with diverse object counts, spatial layouts, and editing instructions, enabling systematic assessment of both semantic alignment and structural controllability. Our main contributions are summarized as follows:
  • $\bullet$
    We introduce QL-Edit, a new task that enables control image editing while ensuring consistency in object count, category, and spatial layout.
  • $\bullet$
    We propose the IMAGHarmony framework, which integrates a HA module and a PNS strategy to enhance structural consistency and generation stability.
  • $\bullet$
    We construct HarmonyBench, a benchmark covering diverse quantities, categories, and layouts, along with metrics for structural and semantic evaluation.
  • $\bullet$
    Extensive experiments show that IMAGHarmony achieves superior performance over state-of-the-art methods in object count accuracy, layout fidelity, and overall generation quality.
2 Related Work
Controllable Image Editing with Diffusion Models. Diffusion models [7, 43, 48, 50, 51, 45, 13] have shown strong potential for high-fidelity and semantically aligned image generation and editing. Prompt-driven editing methods can be broadly categorized into two paradigms: (1) mask-free editing [68, 3, 17] and (2) mask-based editing. Mask-free approaches [3, 17, 20] rely solely on natural language instructions to guide image transformation without explicit spatial constraints. For instance, MagicBrush [68] introduces a large-scale manually annotated dataset for real-world instruction-based editing. InstructPix2Pix [3] trains a conditional diffusion model on synthetic image-instruction pairs generated via GPT-3 [10] and Stable Diffusion, enabling efficient instruction-following. Prompt-to-Prompt [17] and SmartEdit [20] modify prompts during inference to achieve semantically aligned edits without requiring additional training. Mask-based approaches [65, 71, 49] provide spatial control by restricting edits to user-defined or automatically generated regions. LEDITS [57] combines DDPM inversion with semantic guidance to support efficient and flexible real-image editing without architectural modifications. LEDITS++[2] further introduces implicit masking and multi-edit support in a tuning-free, architecture-agnostic setting. InstDiffEdit[71] leverages cross-modal attention to enable instant mask generation without training. Click2Mask [49] simplifies localized editing using point-based mask generation and masked CLIP-based semantic loss. Despite these advances, existing methods [3, 17, 20, 65, 71, 49] usually struggle with multi-object editing, failing to control object count and spatial layout due to weak counting ability and limited structural modeling.
Semantic Guidance with Vision-Language Models. Vision-language models (VLMs) [5, 33, 25] have been increasingly employed to enhance semantic controllability in image editing [44, 52, 70, 40, 16]. Early efforts focus on layout-aware guidance, where LayoutLLM-T2I [47] integrates LLM-driven layout planning with object-aware diffusion to achieve high-fidelity image generation without manual spatial annotations. Similarly, LayoutGPT [9] translates textual descriptions into explicit spatial layouts, improving the model’s comprehension of spatial and numerical semantics. Beyond layout guidance, recent studies explore unified visual-linguistic modeling to advance cross-modal understanding. Emu3 [59] constructs a shared token sequence space that supports multi-turn reasoning and progressive image refinement. UniToken [21] and SEED-X [11] introduce transferable token representations that bridge diffusion inversion with structured language decoding, while Show-o [61] improves semantic consistency via a shared multimodal latent space. In parallel, VLMs have also been leveraged at inference time as scoring functions for semantic alignment and content fidelity. InitNO [14] identifies suboptimal seeds by analyzing attention patterns, whereas LS [63] proposes latent space density as an intrinsic indicator of sample quality. PickScore [23] further simulates human preferences via VLM-based ranking to select semantically aligned outputs without human supervision. However, existing methods [47, 15, 60, 61] typically require extensive training overhead and still struggle to align with user intent or produce stable editing results.
Figure 2: Overview of the IMAGHarmony framework. Given a reference image and a text instruction, we first sample multiple candidate noise seeds and evaluate their semantic alignment using a vision-language model (VLM). The top-$k$ candidates are retained for inference. The harmony-aware attention (HA) module fuses auxiliary textual and visual features to jointly model object count and spatial layout. An image prompt (IP) attention layer injects these layout-aware features into the UNet backbone without modifying its frozen weights. During inference, the preference-guided noise selection (PNS) strategy selects the best candidate seed to guide the full denoising process, leading to edited images that are consistent in both quantity and layout.
3 Method
Quantity-and-layout consistent image editing (QL-Edit) is a novel task that focuses on controllable image manipulation based on textual instructions, while strictly maintaining the original object count and spatial layout from a reference image. Unlike traditional image editing that prioritizes semantic alignment or stylistic fidelity, QL-Edit places equal emphasis on structural integrity, ensuring that modifications do not alter the number or arrangement of key visual elements. To comprehensively evaluate this task, we decompose QL-Edit into the following three representative editing cases:
(1) Class Editing: Alters the object category while preserving the original count and layout (e.g., “Three Apples $\rightarrow$ Three Pineapples”, where both belong to the fruit class, see Figure 1, second row).
(2) Scene Editing: Modifies the background or contextual environment while maintaining object identity, count, and layout (e.g., +“Flowers in Background”, see Figure 1, third row, left).
(3) Style Editing: Applies a stylistic transformation to the image without affecting semantic content or spatial configuration (e.g., +“Cyberpunk Style”, see Figure 1, third row, right).
3.1 Overall Framework
To address the challenge of controlling object counts and spatial layouts in multi-object image editing, we propose IMAGHarmony, a dedicated framework for the QL-Edit task. As illustrated in Figure 2, IMAGHarmony is built upon the denoising UNet architecture of frozen Stable Diffusion XL (SDXL), and incorporates the proposed harmony-aware attention (HA). Additionally, we present a preference-guided noise selection (PNS) strategy at inference time to enhance structural stability and semantic consistency. Specifically, during training, we first generate multiple initial noise seeds and use SDXL to perform denoising conditioned on the given textual prompt, producing a set of candidate images. These candidates are then evaluated by a frozen vision-language model (VLM) for semantic alignment, and the top-$k$ scoring images along with their corresponding seeds are selected for further processing. The auxiliary text and source image are respectively encoded using frozen text and image encoders to extract conditional text and visual features, which are then fed into the HA module (Section  3.2) to obtain harmony feature that explicitly model object count and implicitly capture layout. Subsequently, a learnable IP attention layer is inserted into the "Down4" block of the UNet to inject the harmony feature. To preserve the generative capability of the original model, we additionally apply cross-attention using a “< Null Prompt >” over the frozen UNet. During inference, we adopt the PNS strategy (Section 3.3) to perform a few denoising steps on the candidate seeds, evaluate their semantic consistency, and ultimately select the optimal noise seed to conduct full denoising, resulting in high-quality edited images that align with the desired object count and spatial layout. Finally, we provide a comprehensive summary of the training and inference strategies implemented in IMAGHarmony (Section 3.4).
3.2 Harmony-Aware Attention Module
Motivation. Diffusion models often struggle to accurately capture object count and spatial layout from visual features alone, due to the absence of explicit modeling for quantity and spatial relations. This limitation hinders their effectiveness in structured multi-object editing. To address this, we propose a harmony-aware attention (HA) module that integrates auxiliary textual and visual cues to encode object count and implicitly capture layout. The HA module facilitates more consistent and controllable editing outcomes by enhancing the model’s perception of quantity and structure.
Architecture. As illustrated on the right of Figure 2 (a), the proposed HA module consists of two learnable linear projections and a multi-head cross-attention mechanism, designed to enhance cross-modal alignment of semantic and structural information. Specifically, let $F_{v}\in\mathbb{R}^{B\times N\times D_{v}}$ and $F_{a}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the visual and auxiliary textual features, where $B$ is the batch size, $N$ the number of tokens. $D_{v}$ and $D_{t}$ are the respective feature dimensions. We first project visual feature $F_{v}$ into the auxiliary textual feature space, according to Eq. (1).
$F_{v}^{\prime}=F_{v}W_{\text{proj}},$ (1)
where $F_{v}^{\prime}\in\mathbb{R}^{B\times N\times D_{t}}$ and $W_{\text{proj}}\in\mathbb{R}^{D_{v}\times D_{t}}$ is a learnable matrix. Next, according to Eq. (2), we feed $F_{v}^{\prime}$ as the query and $F_{a}$ as both key and value into a multi-head cross-attention module to capture quantity-and-layout dependencies:
$F_{va}=\text{MultiHead}(F_{v}^{\prime},F_{a},F_{a}),$ (2)
Then, to ensure compatibility with the original visual feature space, we project the fused features back into the visual feature space, as shown in Eq. (3).
$F_{va}^{\prime}=F_{va}W_{\text{align}},$ (3)
where $F_{va}^{\prime}\in\mathbb{R}^{B\times N\times D_{v}}$ and $W_{\text{align}}\in\mathbb{R}^{D_{t}\times D_{v}}$ is a trainable projection. Furthermore, to flexibly control the influence of the injected harmony features and preserve the integrity of the original visual representations, we introduce a scaling hyperparameter $\alpha$. This parameter controls the strength of $F_{va}^{\prime}$ in the final representation $F_{h}\in\mathbb{R}^{B\times N\times D_{v}}$, as defined in Eq. (4).
$F_{h}=\alpha F_{va}^{\prime}+F_{v}.$ (4)
The harmony feature $F_{h}$ encodes object quantity and layout, yielding a coherent and semantically consistent representation for image editing.
Combined with SDXL. To inject harmony feature while preserving the original SDXL architecture, we adopt a dual-branch IP attention. One branch retains the frozen cross-attention module from SDXL, conditioned on a <Null Prompt> to avoid redundant textual interference. In parallel, the second branch incorporates the harmony feature $F_{h}\in\mathbb{R}^{B\times N\times D_{v}}$, which encodes object count and spatial layout. Then, we insert this IP attention module only at the “Down4” block of the UNet, as shown in Figure 2 (a). To maintain consistency with the weight of the pre-train IP-Adapter, the harmony feature $F_{h}$ is first projected through a frozen projection layer originally used for visual conditioning. This avoids introducing trainable parameters in the modality alignment stage and ensures compatibility with the pre-trained space. The projected harmony feature $F_{h}^{{}^{\prime}}$ is then fed into a learnable cross-attention block to interact with the query features. Let $F_{t}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the textual feature, according to Eq. 5, the final attention output is computed.
$Z_{\text{new}}=\text{Softmax}\left(\frac{QK_{t}^{T}}{\sqrt{d}}\right)V_{t}+\beta\cdot\text{Softmax}\left(\frac{QK_{h}^{T}}{\sqrt{d}}\right)V_{h},$ (5)
where $Q=Z_{t}W_{q}$ is the query projected from the noisy latent $Z_{t}$. The text branch uses frozen projections $K_{t}=F_{t}W_{k}^{t}$, $V_{t}=F_{t}W_{v}^{t}$, while the harmony branch uses $K_{h}=F_{h}^{{}^{\prime}}W_{k}^{h}$, $V_{h}=F_{h}^{{}^{\prime}}W_{v}^{h}$ with newly introduced trainable weights. The hyperparameter $\beta$ balances the contribution of the harmony feature. This dual-branch design enables our method to inject structural semantics into the image editing process while preserving the generalization capabilities of the original SDXL model.
3.3 Preference-Guided Noise Selection Strategy
Motivation. Diffusion models are highly sensitive to initial noise [14, 63], especially in multi-object editing where object count and layout must be preserved. Even with the same prompt, different seeds often produce inconsistent results. To address this, we introduce a preference-guided noise selection (PNS) strategy that uses a frozen vision-language model to evaluate low-step generations and select the most semantically aligned seed, improving consistency and stability during inference.
Seed Search. As shown in Figure 2 (a), we perform a seed search during the data preparation stage to pre-select structurally reliable initializations. Specifically, we uniformly sample 100 random seeds $\{s_{i}\}_{i=1}^{100}$, and for each seed $s_{i}$, we generate an image $I_{i}$ using SDXL [46] under a fixed textual prompt. To assess semantic fidelity, we employ CogVLM2 [19] as an evaluator by prompting it with instance-specific questions, such as “How many dogs are in the image?”. The predicted count is compared against the ground-truth, and a binary accuracy score $A_{i}\in\{0,1\}$ is assigned based on correctness. After computing $A_{i}$ for all seeds, we sort the results and select the top-$k$ seeds with the highest semantic alignment to form a candidate set $\mathcal{S}_{\text{top}}$. In practice, we set $k=5$, which balances efficiency and diversity. This seed search strategy ensures that inference begins from semantically aligned noise conditions, reducing layout distortion and object count errors in multi-object editing.
Seed Selection. As illustrated in Figure 2 (b), we utilize the candidate seed set $\mathcal{S}_{\text{top}}$ identified during the seed search phase. For each seed $s_{j}\in\mathcal{S}_{\text{top}}$, we initialize the diffusion model with the candidate noise and perform a shallow denoising process (10 steps) under the given editing instruction (e.g., “Cats”) to generate a candidate image $\hat{I}_{j}$. Each candidate image $\hat{I}_{j}$ is evaluated using CogVLM2 [19], which produces a semantic alignment score $R_{j}\in[0,1]$ reflecting its consistency with the editing instruction. We then identify the best candidate image with the highest alignment score and the best candidate noise. Finally, the best candidate noise is used to reinitialize the model, and a denoising process (30 steps) is performed to obtain the final edited image, which ensures output stability.
3.4 Training and Inference
Training. We optimize only the proposed HA module and IP attention of "Down4" block, while keeping the UNet of SDXL entirely frozen to retain its pre-trained generative capacity. Given the text features $F_{t}$ and the harmony features $F_{h}$, we adopt the standard MSE loss $L_{\text{MSE}}$ [12] defined as:
$L_{\text{MSE}}=\mathbb{E}_{x_{0},\epsilon,F_{t},F_{h},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},F_{t},F_{h},t)\right\|^{2},$ (6)
where $x_{0}$ denotes the ground-truth image, $\epsilon$ is sampled Gaussian noise, $x_{t}$ is the noisy image at timestep $t$, and $\epsilon_{\theta}$ is the predicted noise from the model. The timestep $t\in[0,T]$ is uniformly sampled, with $T$ denoting the total number of denoising steps. This objective guides the model to learn semantic and structural alignment while preserving the frozen generative base backbone.
Inference. During inference, we adopt the proposed PNS strategy to identify the best candidate. Building on the selected candidate, we further incorporate classifier-free guidance (CFG) [18] to enhance conditional generation and ensure consistency in object count and spatial layout. The final denoising prediction is computed as:
$\hat{\epsilon}_{\theta}(x_{t},F_{t},F_{h},t)=w\cdot\epsilon_{\theta}(x_{t},F_{t},F_{h},t)+(1-w)\cdot\epsilon_{\theta}(x_{t},t),$ (7)
where $w$ is the guidance scale used to control the strength of condition.
Figure 3: Examples from HarmonyBench. Each sample includes a source image and a validated count-object caption, verified by YOLO-World [4] and human review.
4 Experiments
Datasets. To fill the gap in benchmarking quantity-and-layout consistent image editing (QL-Edit), we introduce a HarmonyBench dataset for evaluating a model’s ability to preserve object count and spatial layout during editing. The dataset contains 200 image-caption pairs, systematically organized by object count from 1 to 20, with 10 diverse layout examples per count level to ensure broad coverage of numerical and spatial variations. Images are sourced from real-world datasets (COCO [28], OpenImages [24]) and synthesized using SDXL [46], offering rich diversity in visual style, scene composition, and object arrangement. As shown in Figure 3, each caption (e.g., “Two Foxes” and “Twelve Dogs”) is verified through a two-stage pipeline: automatic object counting via YOLO-World [4], followed by manual inspection to ensure count accuracy and object clarity. During evaluation, each reference image is paired with 10 diverse editing prompts covering category changes, scene modifications, and style variations, resulting in a total of 2,000 test instructions.
Metrics. We conducted a comprehensive evaluation of QL-Edit using both objective and subjective metrics. Objective metrics include object count accuracy (OCA) [34], which measures whether the number of generated objects matches the target count; average precision (AP) [6], which quantifies the spatial alignment between generated and reference layouts; and image reward (IR) [62], a learned score that assesses visual quality and instruction alignment. Subjective evaluation includes semantic correctness (SC) [68] and overall quality (OQ) [3], both based on human preference ratings. SC measures whether the generated image faithfully reflects the editing instruction, while OQ captures the overall perceptual quality and user satisfaction. Higher is better for all metrics.
Table 1: Quantitative comparisons on HarmonyBench. IMAGHarmony achieves the best performance.
Method Class Editing ($\uparrow$) Scene Editing ($\uparrow$) Style Editing ($\uparrow$)
OCA AP IR OCA AP IR OCA AP IR
MagicBrush [68] 70.3 69.2 0.437 91.3 90.5 0.104 90.8 76.7 0.216
InstructPix2Pix [3] 66.7 65.3 0.372 91.8 91.1 0.092 89.1 78.2 0.205
LEDITS [57] 45.2 43.1 0.463 77.2 71.9 -0.135 74.4 64.1 -0.253
LEDITS++ [2] 48.4 45.4 0.498 83.5 88.4 -0.187 77.9 63.9 -0.174
IC-Edit [69] 73.2 74.6 0.548 93.7 90.6 0.198 85.5 79.8 0.287
Show-o [61] 35.5 31.0 0.485 80.5 77.5 -0.134 30.6 43.6 0.124
SEED-X [11] 43.1 15.3 0.601 41.5 37.4 0.206 39.2 26.3 0.224
Emu-3 [59] 51.9 33.9 0.585 57.7 66.8 -0.041 45.1 31.2 0.268
IMAGHarmony (Ours) 80.7 75.7 0.693 94.3 91.8 0.217 91.4 90.2 0.315
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Our configuration can be summarized as follows: (1) We initialize the denoising UNet with the pre-trained SDXL  [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, the IP attention with the pre-trained IP-Adapter  [^2] https://huggingface.co/h94/IP-Adapter. (3) The HA module adopts 8 attention heads by default. Input images are resized to $512\times 512$, and a 5% conditional dropout is applied to support classifier-free guidance (CFG). (4) We optimize the model using AdamW with a constant learning rate of $2.5\times 10^{-4}$ and train for 2k steps with a batch size of 1. During inference, we use the DDIM sampler with 30 denoising steps. Unless otherwise specified, the HA module scale, IP attention scale, and CFG scale are set to 1.0, 1.0, and 5.0, respectively.
4.1 Main Results
We quantitatively compare proposed IMAGHarmony with a broad range of state-of-the-art image editing methods across three cases: (1) mask-free editing approaches, including InstructPix2Pix [3] and MagicBrush [68]; (2) mask-based methods, such as LEDITS [57], LEDITS++ [2], and IC-Edit [69]; and (3) VLM-based approaches, including Show-o [61], SEED-X [11], and Emu3 [59].
Figure 4: Qualitative comparisons with several state-of-the-art models on the HarmonyBench dataset.
Quantitative Results. Table 1 summarizes the performance of IMAGHarmony compared to various state-of-the-art (SOTA) methods on the HarmonyBench dataset across three editing tasks: class, scene, and style editing. Our method consistently outperforms all SOTA methods across all metrics, including mask-free, mask-based, and VLM-guided approaches. Specifically, (1) compared to MagicBrush [68], the best-performing mask-free method, IMAGHarmony achieves improvements of 10.4% in OCA and 6.5% in AP on class editing, indicating stronger modeling of object quantity and spatial structure. Similar trends are observed in scene and style editing cases. (2) While the best mask-based method, IC-Edit [69], shows slight improvements over MagicBrush [68] in OCA and AP, benefiting from mask-based spatial priors, it still falls behind IMAGHarmony across all tasks. Notably, our framework achieves superior spatial control without any manual masks, thanks to the proposed HA attention module, which effectively captures quantity and layout representations. (3) Compared to VLM-guided methods, IMAGHarmony surpasses the strongest method, Show-o [61], by 13.8% and 14.3% in OCA and AP on scene editing. More importantly, it achieves significantly higher scores in IR, demonstrating superior semantic alignment. Unlike VLM-based approaches that typically require extensive parameters, IMAGHarmony delivers strong semantic understanding and controllability through its lightweight generative design. Overall, these results highlight the robustness and effectiveness of IMAGHarmony across diverse QL-Edit scenarios.
Qualitative Results. Figure 4 presents qualitative comparisons between IMAGHarmony and several SOTA methods under the QL-Edit setting. (1) In class editing (first three rows), IMAGHarmony accurately modifies object categories (e.g., “Three Apples” → “Three Pineapples”) while faithfully preserving both object count and spatial layout. In contrast, existing methods (e.g., MagicBrush [68], LEDITS [57]) often fail to change the category despite maintaining the correct number of objects. Furthermore, when editing instances with multiple quantities (e.g., “Seven Dogs” → “Seven Bears”), all methods produce incorrect counts and disrupt the original layout, whereas IMAGHarmony remains stable and structurally consistent. (2) In scene editing (rows four–five), our method successfully modifies the background (e.g., + “Fireworks in Background”) while preserving object integrity and spatial position. Other methods (e.g., LEDITS++ [2], IC-Edit [69]) often entangle foreground and background semantics, leading to failed scene modifications. Notably, SEED-X [11] additionally shows inconsistencies in object count. (3) In style editing (last two rows), IMAGHarmony transfers the target style (e.g., “Cyberpunk”) without altering object semantics or disrupting spatial arrangement. In contrast, existing methods frequently over-stylize or distort structure. For example, Show-o [61] mistakenly changes the object category during editing. In summary, IMAGHarmony effectively disentangles content from instructions, delivering semantically faithful and structurally consistent edits, even under complex multi-object.
Figure 5: User study results on HarmonyBench. Our method achieves the best generation result.
User Study. To assess the perceptual quality and semantic fidelity of the edited images, we conducted a user study involving 40 participants. We randomly sampled 50 prompts from the class, scene, and style editing tasks, along with corresponding outputs from all competing methods. Participants were asked to select the top-3 results for each prompt based on two criteria: semantic correctness (SC) [68] and overall quality (OQ) [3]. As shown in Figure 5, our method consistently achieved the highest scores across all metrics, indicating superior semantic alignment, perceptual realism, and user preference. These findings validate the effectiveness of IMAGHarmony from a human-centered evaluation perspective.
4.2 Ablation and Analysis
Effectiveness of the Module. To evaluate the effectiveness of each component in IMAGHarmony, we conduct ablation studies on four variants. Here, B0 removes both auxiliary textual features and our proposed HA module and PNS strategy, injecting source image features via whole IP attention [67]; B1 also excludes auxiliary text and PNS, but injects the source image only through a layout-specific attention layer, similar to InstantStyle [58]; B2 retains PNS but removes the HA module, thus lacking explicit modeling of object quantity and implicit layout reasoning; B3 removes the PNS strategy while keeping HA, using a random seed during inference.
Table 2: Ablation study on HarmonyBench.
Methods OCA ($\uparrow$) AP ($\uparrow$) IR ($\uparrow$)
B0 43.2 21.7 0.161
B1 48.6 56.9 0.193
B2 51.3 49.9 0.367
B3 71.3 68.8 0.322
Ours 88.8 85.9 0.408
As shown in Table 2, B0 performs the worst across all metrics, highlighting the need for structure-aware and semantically guided modeling. B1 improves AP (21.7% → 56.9%) via layout-specific injection but still lags in OCA and IR, suggesting limited gains from shallow spatial cues. B2 introduces PNS, yielding better OCA and IR, yet lacks explicit modeling of object quantity and layout. B3 benefits from the HA module, achieving higher OCA and AP , but suffers in IR due to unstable seed initialization. In contrast, the full IMAGHarmony model with both HA and PNS, achieves the best performance, confirming the complementary benefits of harmony features and seed-aware inference.
Figure 6: Attention map comparison. (a) w/o HA. (b) w/ HA. HA module better captures object count and layout.
Visualization of the Feature Map. To further validate the effectiveness of the proposed HA module in modeling object quantity and spatial layout, we visualize and compare attention distributions with and without the HA module. From Figure 6, the heatmaps are generated by averaging activation maps from the "Down4" blocks of the UNet. The source image contains multiple well-distributed objects, representing a typical multi-object editing scenario. In Figure 6 (a), without the HA module (w/o HA) exhibits dispersed attention with weak structural focus, often leading to missing or misaligned objects. In contrast, Figure 6 (b) shows that our model with HA produces sharper, more concentrated attention aligned with the object regions. These visualizations demonstrate that the HA module effectively enhances the model’s ability to capture object count and spatial structure, thereby improving semantic and structural consistency in QL-Edit.
Figure 7: Performance comparison of injection location. Injecting at D4 yields the highest accuracy in both OCA and AP.
Injection Location Analysis. As shown in Figure 7, we evaluate the impact of injecting the harmony feature at different locations within the UNet, including downsampling (D), middle (M), and upsampling (U) blocks. D1 refers to the first downsampling block. The results demonstrate that inserting the harmony feature at the fourth downsampling block (D4), yields the highest performance across both object count accuracy (OCA) and average precision (AP), significantly outperforming all other configurations. This suggests that introducing structure-aware guidance at D4 provides a better balance between semantic and spatial structure, allowing the model to capture quantity and layout consistency more effectively. In contrast, injecting at shallower or deeper layers fails to model structural relationships, resulting in inferior performance.
Figure 8: Effect of attention weights on editing quality. Setting both $\alpha$ and $\beta$ to 1.0 produces the best results.
Impact of Hyperparameters. We further study the impact of two key hyperparameters, $\alpha$ and $\beta$, which control the influence of the HA module and IP Attention modules, respectively, as defined in Eq. (4) and Eq. (5). We vary each parameter independently from 0.0 to 1.0 in increments of 0.2, while keeping the other fixed at 1.0. As shown in Figure 8, the best editing quality occurs when both $\alpha=1.0$ and $\beta=1.0$. In contrast, lower values lead to unstable object quantity and spatial layout, resulting in reduced structural fidelity and semantic consistency.
5 Conclusion
In this paper, we introduce QL-Edit, a new task that emphasizes quantity and layout consistency in multi-object image editing. To tackle the challenges of structural control and generation stability, we propose IMAGHarmony, a consistency-aware framework that combines a harmony-aware attention mechanism with a preference-guided noise selection strategy. Our approach effectively aligns object semantics and spatial structures, outperforming existing methods across various metrics. We also release HarmonyBench, a comprehensive benchmark for quantity-layout controllable editing.